OpenStudy (anonymous):

A retailer has determined that the cost C of ordering and storing x units of a product is c=6x+900000/x (a) Write the expression for cost as a single fraction. (b) Determine the cost for ordering and storing - x=240 units of this product.

OpenStudy (anonymous):

\[Cost = 6x + \frac{900000}{x} = \frac{6x^2 + 900000}{x}\] Plug in the value of x = 240 units \[Cost = \frac{6 \times 240^2 + 900000}{240} = \frac{6 \times 57600 + 900000}{240}\]

OpenStudy (king):

5190